OpenStudy (anonymous):

Kaihi can eat 32 lollies in one hour. Her sister can eat the same amount in 3 hours Howlong will it take both girls to eat 32 lollies? Can you tell mehow to do it easily no straight up answers

OpenStudy (anonymous):

lollies. lol

OpenStudy (anonymous):

you there? @CassiusDelaney

OpenStudy (anonymous):

Yerrr i am now...

OpenStudy (anonymous):

So can anyone help??

OpenStudy (anonymous):

yeh sry...

OpenStudy (anonymous):

so rate * time... is amount...

OpenStudy (anonymous):

32 lollies (lol) per hour for kaihi and 32 per 3 hours for her sister...

OpenStudy (anonymous):

Forget about the 32 lollies and consider it as one task to accomplish. Then to accomplish that task Kaihi (k) spent k hours and her sister (s) s hour. To accomplish the task together, let say, they spent x hours, Thus this must hold true:\[\frac{ 1 }{ k }+\frac{ 1 }{ s }=\frac{ 1 }{ x }\]\[\frac{ 1 }{ 1 }+\frac{ 1 }{ 3 }=\frac{ 1 }{ x }\] soving for x you will get \[x =\frac{ 3 }{ 4 }hr\]

OpenStudy (anonymous):

so 32*t + 32*t/3 = 32 lollies

OpenStudy (anonymous):

Did you get it?

OpenStudy (anonymous):

nahhh sorry is there a easyer way or the fractions work just need to no how you found the answers

OpenStudy (anonymous):

1/k of the task is completed by K 1/s of the task completed by S 1/x of the task completed by both.

OpenStudy (anonymous):

how did you get the denominator?

OpenStudy (anonymous):

Maybe a little simpler: Kaihi 96 in 3 hours Sister 32 in 3 hours Together 128 in 3hours and now you want That's for 128 lollies and u want 32 so 3 hours * 32/128 = 96/128 = 3/4 hour